# Spins of Black Holes in X-ray Binaries and the Tension with the Gravitational Wave Measurements

### v2

@@ Acknowledgements
We thank Sudeb Datta, Shane Davis, Gulab Dewangan, Julian Krolik, Piero Madau, Joanna Mikołajewska, Ranjeev Misra, Andrew Mummery, Andrzej Niedźwiecki, and Thomas Tauris for their comments. We also thank the two referees for their thoughtful suggestions. We acknowledge support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138. Alexandra Olejak acknowledges funding from the Netherlands Organization for Scientific Research (NWO) as part of the Vidi research program BinWaves (project number 639.042.728, PI: de Mink). Grégoire Marcel and Alexandra Veledina acknowledge support from the Academy of Finland grant 355672. Debora Lančová acknowledge the Czech Science Foundation (GAČR) project No. 25-16928O. Nordita is supported in part by NordForsk.

### v3

@@ Acknowledgements
We thank Sudeb Datta, Shane Davis, Gulab Dewangan, Julian Krolik, Piero Madau, Joanna Mikołajewska, Ranjeev Misra, Andrew Mummery, Andrzej Niedźwiecki, and Thomas Tauris for their comments. We also thank the two referees for their thoughtful suggestions. We acknowledge support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138. Alexandra Olejak acknowledges funding from the Netherlands Organization for Scientific Research (NWO) as part of the Vidi research program BinWaves (project number 639.042.728, PI: de Mink). Grégoire Marcel and Alexandra Veledina acknowledge support from the Academy of Finland grant 355672. Nordita is supported in part by NordForsk. Debora Lančová acknowledges the Czech Science Foundation (GAČR) project No. 25-16928O.